# Recitation over Reasoning: How Cutting-Edge Language Models Can Fail on Elementary School-Level Reasoning Problems?

### v1

@@ 2 Related Work
LLM robustness. While LLM achieves tremendous success, there has been persisting concerns about the limited robustness of LLMs [94, 76]. For example, LLMs have been well known for making mistakes in comparing $9.8$ and $9.11$ [76] and counting “r”s in “strawberry” [77]; there have also been many works that question LLM’s robustness when confronted with out-of-distribution data [56, 85], incorrect/incomplete commands [79, 90], complex calculations [94], symbolic relations [46], and order of choices in multiple choice questions [91]. The work with the most similar findings to ours is done by Qiu et al. [54], who reports similar vulnerability of OpenAI-o1 [30] to subtly changed key conditions as spurious correlation to often-asked concepts, and proposed to fix the issue by using symbolic graphics data. However, their work and proposed solution focus on visual understanding of multi-modal LLMs, while we find that the recitation of solution templates are even more severe in pure text-based problems.
Multi-modal LLMs. As the inherent limit of languages [25] and corpus depletion [67] quickly becomes a major obstacle for AGI, researchers quickly turn to other modalities, such as vision [13] and speech/audio [37, 18] for extra input sources. As humans take the most information from vision [27], Vision Language Models (VLMs) such as OpenFlamingo [6], Llava [39, 40], Qwen-VL [8, 9] and GPT-4v/-4o [47, 26] have become the prevailing paradigm for multimodal LLMs, and made unique progress on multiple areas beyond LLMs, such as robotics [70, 16] and autonomous driving [64, 78, 84]. VLMs are also evaluated by part of our benchmark, and they exhibit the same recitation problem. Moreover, researchers have discovered the spurious correlation phenomenon [66, 24] by VLMs; such correlation between often-tested notions (e.g. famous optical illusions) and modified inputs [54] becomes part of the source for improper recitation in our benchmark.
@@ 3 RoR-Bench
Figure 2: Examples of problems in our benchmark; for better readability, we marked the modified part red. Despite that we build a Chinese benchmark, OpenAI-o1-1217 [30] also fails with our English translation for these examples.
@@ 5 Discussion and Conclusion
Limitations and Future Works. Currently, our benchmark is Chinese-only due to the language limitation of human annotators and moderators, which may cause an edge on performance for LLMs by Chinese companies such as Ernie-4.5 and Hunyuan Turbo-S (note the main message, significant performance decrease after modification, is not affected). Though our message to convey is already strong with the current results (and preliminary English translation tests in this paper suggest that LLMs will other struggle on the other languages), to expand such benchmark to multiple languages will be an important future work. A more important and fundamental avenue for future research is to find an effective way for LLMs to overcome such problem of recitation over reasoning without relying too much on user’s clarifications or being too harsh on typos.
@@ Acknowledgements
We would like to thank the following annotators for curating this dataset: Zhihui Song, Xinyue Zhang, Leyao Li, Zhuowei Li, Li Jiang, Shaofeng Pan, Chenxi Wang, Lin Wang, Bei Wang, Xin Luo, Jiahe Zhao, Mengyao Yan, Yutong Yan, Jiaojiao Chen, Tiantian Chen, Yi Chen and Yuxiao Han, and the moderators: Xiaowen Guo, Leyao Li, Shaofeng Pan, Chenxi Wang, Lin Wang and Jiahe Zhao. We would also like to thank Tong Sun, Zezhong Ma, Zaiyuan Wang, Aowei Ji and Qinyan Zhang for their contribution and discussion to the idea.

### v2

@@ 2 Related Work
LLM robustness. While LLM achieves tremendous success, there has been persisting concerns about the limited robustness of LLMs [98, 79]. For example, LLMs have been well known for making mistakes in comparing $9.8$ and $9.11$ [79] and counting “r”s in “strawberry” [80]; there have also been many works that question LLM’s robustness when confronted with out-of-distribution data [58, 88], incorrect/incomplete commands [82, 93], complex calculations [98], symbolic relations [48], and order of choices in multiple choice questions [95]. Recently, the vulnerability of LLM reasoning under perturbed conditions has attracted the researcher’s attention, for example, LLM’s math ability under conditions with irrelevant context [60] or extended reasoning steps [99]. The most similar works to ours are done by Zhao et al. [94] and Huang et al. [25], both of which include math problems with subtly but fundamentally changed conditions. However, both works do not contain multi-modal problems, and their original problems without trap contains only math problems with more complex knowledge (e.g. number theory or precalculus). On the contrary, our benchmark contains more reasoning problems with less prior knowledge, and shows larger gap between original and modified problems.
Multi-modal LLMs. As the inherent limit of languages [27] and corpus depletion [70] quickly becomes a major obstacle for AGI, researchers quickly turn to other modalities, such as vision [13] and speech/audio [39, 18] for extra input sources. As humans take the most information from vision [29], Vision Language Models (VLMs) such as OpenFlamingo [6], Llava [41, 42], Qwen-VL [8, 9] and GPT-4v/-4o [49, 28] have become the prevailing paradigm for multimodal LLMs, and made unique progress on multiple areas beyond LLMs, such as robotics [73, 16] and autonomous driving [67, 81, 87]. VLMs are also evaluated by part of our benchmark, and they exhibit the same recitation problem. There are some recent works that provide explanations for such issue. For example, some argue that the problem comes from spurious correlation [69, 24], where correlation between often-tested notions (e.g. famous optical illusions) and modified inputs becomes part of the source for improper recitation, and reports similar issues to our findings [56]; others argue that the problem comes from inefficient decoding [26] or memorization [100], the latter of which resembles our argument.
@@ 3 RoR-Bench
Figure 2: Examples of problems in our benchmark; for better readability, we marked the modified part red. Despite that we build a Chinese benchmark, OpenAI-o1-1217 [32] also fails with our English translation for these examples. See Appendix B.6 for more experiments on the English translation.
@@ 5 Discussion and Conclusion
Limitations and Future Works. Currently, our benchmark is Chinese-only due to the language limitation of human annotators and moderators, which may cause an edge on performance for LLMs by Chinese companies such as Ernie-4.5 and Hunyuan Turbo-S (note the main message, significant performance decrease after modification, is not affected). Though our message to convey is already strong with the current results (and preliminary English translation tests in this paper suggest that LLMs will other struggle on the other languages), to expand such benchmark to multiple languages will be an important future work. A more important and fundamental avenue for future research is to find an effective way for LLMs to overcome the problem of recitation over reasoning without over-reliance on user’s clarifications or being too harsh on typos.
@@ Acknowledgements
We would like to thank the following contributors of this dataset: Zhihui Song, Xinyue Zhang, Leyao Li, Zhuowei Li, Li Jiang, Shaofeng Pan, Chenxi Wang, Lin Wang, Bei Wang, Xin Luo, Jiahe Zhao, Mengyao Yan, Yutong Yan, Jiaojiao Chen, Tiantian Chen, Yi Chen and Yuxiao Han, and the following people for double-checking our dataset: Xiaowen Guo, Leyao Li, Shaofeng Pan, Chenxi Wang, Lin Wang and Jiahe Zhao. We would also like to thank Tong Sun, Zezhong Ma, Zaiyuan Wang, Aowei Ji and Qinyan Zhang for their contribution and discussion to the idea.
@@ Appendix A Prompts
See https://chatgpt.com/share/67edc40f-e658-800d-9bcd-7cb268a6f8c9 for another example of using the word “floating” and https://chatgpt.com/share/67ee1ba2-089c-800d-ad01-d7ee50dab39d for “drifting”. Interestingly, o1 can recognize the issue when we use the two words together (https://chatgpt.com/share/67ee1bda-5060-800d-ad44-b6c4e5741185), but not separately.
@@ B.5 Failure Example on Original Problem with “Forced Correct” Prompt
B.6 English Version of Fig. 2
The response for OpenAI-o1-1217 for our English translation of the text problems in Fig. 2 can be seen in https://chatgpt.com/share/67f45d45-a694-800d-a4c8-3c763e93400f. In later experiments, the response of GPT-4.5 points out that there is a possibility of ambiguity - “traveling in opposite directions” could also possibly mean traveling towards each other, as if the two cars are located in the west and east respectively, the car on the west will need to travel east and the car on the east will need to travel west (https://chatgpt.com/share/67f45e46-5eb0-800d-8f1c-5dec1f3c9f1d), which are indeed opposite directions. To further address such concerns, we test the result of literal translation from Chinese (“traveling back to back”), and the following version with much more explicit explanation:
Two cars start simultaneously from two cities that are 300 km apart, one on the east and the other on the west. The car on the east travels towards the east, and the car on the west travels towards the west. One car has a speed of 60 km/h, while the other has a speed of 70 km/h. How many hours will it take for them to meet?
Surprisingly, we found OpenAI-o1-1217 still fails on both the literal translation (https://chatgpt.com/share/67f4601a-149c-800d-8be1-d4ebe97c315b) and the explicit version above (https://chatgpt.com/share/67f444ed-8148-800d-810d-62c54ea1636a).